I need to find the rule of the infinite series starting with
1, 2, 4, 8,...
i tried multiplying, using powers, adding and subtracting..
and nothing's worked! please help and give me tips to do it
faster.. thanks:)

Answer by stanbon(75887)

I need to find the rule of the infinite series starting with
1, 2, 4, 8,...
-----------
1st: 1 = 2^0
2nd: 2 = 2^1
3rd: 4 = 2^2
4th: 8 = 2^3
.....
nth: 2^(n-1)
------
Rule a(n) = 2^(n-1)
===========================
Cheers,
Stan H.
===========================